Using the method of undetermined coefficients, write the trial solution of the equation d^2y/dx^2+2dy/dx+5y=x e^(–1) cos2x and hence solve it.
1
Expert's answer
2020-03-02T09:47:22-0500

"y''+2y'+5y=e^{-x}\cos 2x"

Characteristic equation: "\lambda^2+2\lambda+5=0", which has roots "\lambda_1=-1-2i" and "\lambda_2=-1+2i"

Trial solution is "y=Axe^{-x}\cos 2x+Bxe^{-x}\sin 2x"

"y'=Ae^{-x}\cos 2x-Axe^{-x}\cos 2x-2Axe^{-x}\sin 2x+"

"+Be^{-x}\sin 2x-Bxe^{-x}\sin 2x+2Bxe^{-x}\cos 2x"

"y''=Axe^{-x}\cos 2x-4Axe^{-x}\cos 2x-2Ae^{-x}\cos 2x-"

"-4Ae^{-x}\sin 2x+4Axe^{-x}\sin 2x+Bxe^{-x}\sin 2x-"

"-4Bxe^{-x}\sin 2x-2Be^{-x}\sin 2x+4Be^{-x}\cos 2x-"

"-4Bxe^{-x}\cos 2x"

"e^{-x}\cos 2x=y''+2y'+5y=4Be^{-x}\cos 2x-4Ae^{-x}\sin 2x"

"B=\frac{1}{4}, A=0"

Answer:

Trial solution is "\frac{1}{4}xe^{-x}\sin 2x"

General solution is "Ce^{-x}\cos 2x+De^{-x}\sin 2x+\frac{1}{4}xe^{-x}\sin 2x"